OpenStudy (anonymous):

Hello people! Please Please can someone help me??? I really need some help with this question cause I dont understand anything... okay so the question is: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. x^2 - 36 = 0 and x^2 = 8x - 12 Oh and please show work so I can understand whats going on...

OpenStudy (anonymous):

x^2 - 36=0 or x^2 -6^2=0 or (x+6)(x-6)=0 or x=+6 0r -6

OpenStudy (anonymous):

x^2 = 8x - 12 x^2 -8x + 12=0 or x^2 -6x -2x + 12=0 or x(x-6) -2(x-6) =0 or (x-6)(x-2)=0 or x=6 or x=2 hence from both the sol. of equation we have x=6

OpenStudy (anonymous):

but wait could you explain the first one a little better? The second is perfect ( Thank you!) but I still dont see how you got the first questions answers.....

OpenStudy (anonymous):

On the first one you could add 36 to both sides: x^2=36. Take the square root of each side and x=6 or -6.

OpenStudy (anonymous):

oh ok i think i get it :) thanks!